# Generative Adversarial Networks for High-Dimensional Item Factor Analysis: A Deep Adversarial Learning Algorithm

### v2

@@ 2.5 Our Contributions to the Literature
Given that VAEs have been a powerful approach to perform exploratory IFA (Urban and Bauer,, 2021), the innovative contribution of the AVB method is to bring richer inference model choices for the implementation by deep learning. Over the past decade, deep learning techniques have thrived (LeCun et al.,, 2015), and IFA has derived considerable benefits from these advances. When applying neural networks to estimate parameters in IFA, key distinctions arise in the design of the network architecture and the choice of loss function. A straightforward strategy is to feed one-hot encodings [^1] In one-hot encoding, each respondent or item is represented by a binary vector whose length matches the total number of respondents or items. Exactly one element of this vector is set to 1 (indicating the specific respondent or item), and all other elements are set to 0. for both respondents and items into two separate feedforward neural networks: one infers the latent variables of the respondents, while the other estimates the item parameters. This approach typically employs a variation of the Joint Maximum Likelihood (JML) for its loss function, relying on the product of conditional probabilities in Equation (3) across all respondents (Tsutsumi et al.,, 2021; Zhang and Chen,, 2024). However, as demonstrated by Ghosh, (1995), the JML solution leads to inconsistent estimates for item parameters.
@@
Building on insights from Sections 2.3 and 2.4, we observe that the major contrast between AVB and conventional VAE approaches centers on the inference model. While generative models demand precise techniques for latent-variable estimation, basic Gaussian assumptions might restrict the range of distributions in standard VAEs. Although researchers have attempted to design more sophisticated neural network architectures for VAE inference, these solutions may still lack the adaptability of the black-box method employed by AVB (Mescheder et al.,, 2017).
To address this limitation, the AVB framework modifies $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ from Equation (10) into a fully black-box inference mechanism that cooperates with adversarial objectives. Through this process, the algorithm derives item parameter estimates $\bm{\theta}^{*}$ that optimizes the likelihood of the observed data, while also constructing a distribution $q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})$ that closely approximates the true posterior $p_{\bm{\theta}^{*}}(\bm{z}\mid\bm{x})$.
From a theoretical perspective, Mescheder et al., (2017) established that AVB can achieve performance on par with at least that of VAEs in terms of likelihood, as evidenced by the following theorem:
Theorem 2.2.
Assume that $T$ can represent any function of $\bm{x}$ and $\bm{z}$. If $(\bm{\theta}^{*},\bm{\phi}^{*},T^{*})$ defines a Nash-equilibrium [^2] Nash equilibrium is a central concept in game theory wherein each player’s strategy is optimal, given the strategies of all other players. In the context of GANs, this equilibrium arises in the two-player adversarial optimization: once the generator and the discriminator reach a Nash equilibrium, neither can unilaterally adjust its parameters to further improve its objective without a corresponding change by the other. of the two-player game defined by the Equations (11) and (13), then
$T^{*}(\bm{x},\bm{z})=\log q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})-\log p(\bm{z})$ (14)
and $(\bm{\theta}^{*},\bm{\phi}^{*})$ is a global optimum of the variational lower bound in the Equation (9).
@@
Figure LABEL:fig:comp_avb_vae visually contrasts these two approaches. Rather than a standard approach where random noise is added only at the final step, AVB incorporates $\bm{\epsilon}$ as an additional input earlier in the inference process. The distribution of latent variables in VAE is assumed to be based on normal distributions, but AVB is free from this assumption with richer probabilistic model choice. As neural networks can almost represent any probability density in the latent space (Cybenko,, 1989), this design empowers the network to represent intricate probability distributions more effectively, thus expanding the scope of patterns it can capture.
@@
Under misspecification such as inappropriate assumptions of latent variable distributions, AVB’s architecture can deliver a higher marginal likelihood theoretically. Meanwhile, improving the marginal log-likelihood can reduce the reconstruction error $-\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]$, which measures the disparity between real and reconstructed data, and can be reduced by increasing marginal likelihood according to Equation (8). In the context of IRT, AVB thus stands out by capturing a more accurate respondents’ latent traits. Research in the computer science field also highlights that the combination of GAN and VAE can extract more discriminative information and provide more precise and clear model in the latent variable space (e.g., Makhzani et al.,, 2015; Hou et al.,, 2019).
@@ 3.1 Neural Network Design
The Algorithm 1 can be represented by the schematic diagram in the Figure 5. The encoder $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ and discriminator $T(\bm{x},\bm{z})$ are designed to be multiple-layer feedforward neural networks (FNNs) in this study. A FNN is a type of artificial neural network where information flows in one direction, from the input layer to the output layer, through one or more hidden layers, without forming any cycles or loops. It is the simplest form of a neural network and by increasing the number of neurons in the hidden layer, any continuous function on a closed interval can be approximated at an arbitrary precision (Cybenko,, 1989).
Figure 5: Schematic diagram of the Adversarial Variational Bayes theorem.
FNN consists of layers of neurons, where each layer $l$ is represented mathematically by the following recursive relation: given the weight matrix $\bm{W}^{(l)}$ connecting layer $l-1$ and $l$, the $P_{l-1}\times 1$ output vector $\bm{h}^{(l-1)}$ from the previous layer $l-1$ and the $P_{l}\times 1$ bias vector $\bm{b}^{(l)}$ for layer $l$, the $P_{l}\times 1$ output $\bm{h}^{(l)}$ of the neurons in layer $l$ is defined by
@@
Parameter initialization is a critical step in training neural networks, as it sets the initial values of weights and biases before optimization begins. Proper initialization prevents issues like vanishing or exploding gradients, which can hinder learning and slow convergence. For encoder network and discriminator network, Kaiming initialization (He et al.,, 2015) is chosen for the weight and bias since it is tailored for asymmetric activation function like GELU. As GELU almost drops most negative activations and reduces variance of parameters in subsequent layers, Kaiming initialization preserves variance across layers. Considering $P_{l}\times P_{l-1}$ matrix $\bm{W}_{0}^{(l)}=[w_{p_{l},p_{l-1},0}]$ and $P_{l}\times 1$ vector $\bm{b}_{0}^{(l)}=(b_{p_{l},0})$, Kaiming initialization is defined as follows:
where $\mathcal{U}(a,b)$ represents the uniform distribution with lower bound $a$ and upper bound $b$. For the decoder, which can be regarded as a one-layer neural network with sigmoid function, Xavier initialization is chosen for the loading and intercept parameters. Xavier initialization (Glorot and Bengio,, 2010) is designed for symmetric activations (e.g., sigmoid) and balances the variance of activations and gradients by considering both the number of input and output neurons. For item $j=1,\dots,M$, the $P\times 1$ vector of loadings $\bm{\beta}_{j,0}=(\beta_{p,j,0})$ and the $(C_{j}-1)\times 1$ vector of intercepts $\bm{\alpha}_{j,0}=(\alpha_{k,j,0})$ are initialized by: for $p=1,\dots,P$ and $k=1,\dots,C_{j}-1$,
In FNNs, parameter optimization is often performed using Stochastic Gradient Descent (SGD) methods, which iteratively adjust the model’s weights and biases to minimize a chosen loss function (LeCun et al.,, 2015). SGD is a foundational method where the model parameters are updated by computing the gradient of the loss with respect to the parameters for a randomly sampled batch of training data. Specifically, the update rule is given by $\bm{\omega}_{t+1}=\bm{\omega}_{t}-\eta\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega})$, where $\bm{\omega}_{t}$ represents the parameters (e.g., $\bm{\theta},\bm{\phi},\bm{\psi}$ in the Algorithm 1) at step $t$, $\eta$ is the learning rate, and $\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega})$ is the gradient. Modern extensions of SGD methods, such as AdamW (Loshchilov,, 2017), improve on SGD by adaptive learning rates and decoupled weight decay. AdamW separates the weight decay term from the gradient updates, which enables better generalization in deep learning models because it explicitly penalizes large weights through a regularization term without conflating it with the optimization step. The parameter update rule in AdamW includes momentum and second-order moment estimation, and is expressed as follows:
[…]
where $\bm{m}_{t}$ and $\bm{v}_{t}$ are the first and second moment estimates, $\beta_{1}$ and $\beta_{2}$ are exponential decay rates for the first and second moment estimates, $\lambda$ is the weight decay coefficient, and $\epsilon$ is a small constant for numerical stability. The hyperparameters are set to the default values ($\beta_{1}=0.9,\beta_{2}=0.999,\lambda=0.01,\epsilon=1\times 10^{-8}$) in this study, as they are practically recommended (Loshchilov,, 2017; Reddi et al.,, 2019). AdamW has become a popular choice in optimizing neural networks due to its ability to converge faster and improve model performance by effectively mitigating overfitting caused by large parameter values (Zhuang et al.,, 2022).
@@
In the Algorithm 1, the gradient descent step tries to force $T(\bm{x},\bm{z})$ close to $T^{*}(\bm{x},\bm{z})$ but the gap might fail to be sufficiently small. In traditional AVB, the adversary tries to distinguish between samples from the true posterior distribution and those from the variational approximation, $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$. However, this task can be quite challenging if there’s a large difference between $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the prior distribution $p(\bm{z})$.
@@ 3.2 Adaptive Contrast Technique
To closely approximate the $q_{\phi}(\bm{z}\mid\bm{x})$, a tractable density choice of $r_{\alpha}(\bm{z}\mid\bm{x})$ is a Gaussian distribution with a diagonal covariance matrix, which matching the mean $\bm{\mu}(\bm{x})$ and variance $\bm{\sigma}^{2}(\bm{x})$ of $q_{\phi}(\bm{z}\mid\bm{x})$. Therefore, $r_{\alpha}(\bm{z}\mid\bm{x})$ is adaptive according to updating $\phi$. The KL term can also be approximated by the Theorem 2.1 and $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),r_{\alpha}(\bm{z}\mid\bm{x})\right]$ can be much smaller than $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\theta}(\bm{z})\right]$ so it is much easier to construct an effective discriminator. Moreover, reparameterization trick can further simplify the KL term into
$\text{KL}\left[\tilde{q}_{\bm{\phi}}(\tilde{\bm{z}}\mid\bm{x}),r_{0}(\tilde{\bm{z}})\right],r_{0}(\tilde{\bm{z}})\sim\mathcal{N}(\bm{0},\bm{I}),\tilde{\bm{z}}=\frac{\bm{z}-\bm{\mu}(\bm{x})}{\bm{\sigma}(\bm{x})}.$ (23)
@@ 3.3 Importance-weighted Technique
If the number of importance-weighted (IW) samples $R=1$, the IW-ELBO is reduced to the ELBO. As $R$ increases, IW-ELBO becomes more similar to the marginal likelihood than ELBO (Burda et al.,, 2015). Therefore, optimizing IW-ELBO, which is the importance-weighted Objective (9) or (22), achieves a better approximation to the MML estimator with slightly low computational efficiency due to more samples but GPU computation source can ensure the computation time under control and even less than the time required by traditional CPU computation.
@@ 3.4 Implementation Summary
                   $w_{i,r,s}\leftarrow\exp\left[-{T}_{\bm{\psi}}\left(\bm{x}_{i},\tilde{\bm{z}}_{i,r,s}\right)+\frac{1}{2}||\tilde{\bm{z}}_{i,r,s}||^{2}+\log p_{\bm{\theta}}\left(\bm{x}_{i},\bm{z}_{i,r,s}\right)\right]$;
@@
[…]
@@
       Optimization step: Perform SGD-updates via AdamW: for $\bm{g}_{t}=(\bm{g}_{\bm{\theta}_{t}}^{T},\bm{g}_{\bm{\phi}_{t}}^{T},\bm{g}_{\bm{\psi}_{t}}^{T})$;
       $\tilde{\bm{g}}_{t}=\bm{g}_{t}+\lambda(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})$, $\bm{m}_{t+1}=\beta_{1}\bm{m}_{t}+(1-\beta_{1})\tilde{\bm{g}}_{t}$, $\bm{v}_{t+1}=\beta_{2}\bm{v}_{t}+(1-\beta_{2})\tilde{\bm{g}}_{t}^{2}$;
       $\hat{\bm{m}}_{t+1}=\bm{m}_{t+1}/(1-\beta_{1}^{t+1})$, $\hat{\bm{v}}_{t+1}=\bm{v}_{t+1}/(1-\beta_{2}^{t+1})$;
       $(\bm{\theta}_{t+1}^{T},\bm{\phi}_{t+1}^{T},\bm{\psi}_{t+1}^{T})=(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})-\eta\frac{\hat{\bm{m}}_{t+1}}{\sqrt{\hat{\bm{v}}_{t+1}}+\epsilon}-\eta\lambda(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})$;
@@
Determining the optimal number and size of hidden layers in neural network models is essential for high performance. However, there is no panacea solution, as the architecture often depends on the specific application and the characteristics of the data. Empirical studies have shown that one or two hidden layers are typically sufficient to represent most continuous functions efficiently (Huang and Babri,, 1997; Huang,, 2003), and a maximum of $12n+$ 1 nodes can capture complex patterns (Hecht-Nielsen,, 1987). This study conducted several experiments across different tasks and selected the hyperparameter values that yielded the best performance from the ranges of $\{1,2\}$ for the number of hidden layers and $\{64,128,256,512\}$ for the number of neurons. The experimental results indicated that a more powerful discriminator, which is achieved by increasing its depth or width, enhances feature learning and, consequently, improves generator performance.

### v3

@@ 2.5 Our Contributions to the Literature
Given that VAEs have been a powerful approach to perform exploratory IFA (Urban and Bauer,, 2021), the innovative contribution of the AVB method is to bring richer inference model choices for the implementation by deep learning. Over the past decade, deep learning techniques have thrived (LeCun et al.,, 2015), and IFA has derived considerable benefits from these advances. When applying neural networks to estimate parameters in IFA, key distinctions arise in the design of the network architecture and the choice of loss function. A straightforward strategy is to feed one-hot encodings [^3] In one‐hot encoding, each respondent or item is represented by a binary vector whose length matches the total number of respondents or items. Exactly one element of this vector is set to 1 (indicating the specific respondent or item), and all other elements are set to 0. for both respondents and items into two separate feedforward neural networks: one infers the latent variables of the respondents, while the other estimates the item parameters. This approach typically employs a variation of the Joint Maximum Likelihood (JML) for its loss function, relying on the product of conditional probabilities in Equation (3) across all respondents (Tsutsumi et al.,, 2021; Zhang and Chen,, 2024). However, as demonstrated by Ghosh, (1995), the JML solution leads to inconsistent estimates for item parameters.
@@
Building on insights from Sections 2.3 and 2.4, we observe that the major contrast between AVB and conventional VAE approaches centers on the inference model. Figure 5(a) visually contrasts these two approaches. While generative models demand precise techniques for latent-variable estimation, basic Gaussian assumptions might restrict the range of distributions in standard VAEs. Although researchers have attempted to design more sophisticated neural network architectures for VAE inference, these solutions may still lack the expressiveness of the black-box method employed by AVB (Mescheder et al.,, 2017).
@@
To address this limitation, the AVB framework modifies $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ into a fully black-box neural network $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$. Rather than a standard approach where random noise is added only at the final step, AVB incorporates $\bm{\epsilon}$ as an additional input earlier in the inference process. As neural networks can almost represent any probability density in the latent space (Cybenko,, 1989), this design empowers the network to represent intricate probability distributions free from the Gaussian assumption, thus expanding the scope of patterns it can capture. However, such implicit $q_{\bm{\phi}}$ lacks a tractable density, $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ cannot be directly computed so another discriminator neural network learns to recover $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ via adversarial process. The recovery of $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ imposes a prior, such as a standard normal, on $\bm{z}$. This prior anchors the location and scale of $\bm{z}$, thereby alleviating the translational and scaling indeterminacies that cause model non-identifiability (De Ayala,, 2013).
(a) Schematic comparison of the encoder and decoder designs for the AVB method and a standard VAE.
Through the adversarial process, the algorithm derives item parameter estimates $\bm{\theta}^{*}$ that optimizes the likelihood of the observed data, while also constructing a distribution $q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})$ that closely approximates the true posterior $p_{\bm{\theta}^{*}}(\bm{z}\mid\bm{x})$. From a theoretical perspective, Mescheder et al., (2017) established that AVB can achieve performance on par with at least that of VAEs in terms of likelihood, as evidenced by the following theorem:
Theorem 2.2 (Mescheder et al.,, 2017).
Assume that $T$ can represent any function of $\bm{x}$ and $\bm{z}$. If $(\bm{\theta}^{*},\bm{\phi}^{*},T^{*})$ defines a Nash-equilibrium [^4] Nash equilibrium is a central concept in game theory wherein each player’s strategy is optimal, given the strategies of all other players. In the context of GANs, this equilibrium arises in the two‐player adversarial optimization: once the generator and the discriminator reach a Nash equilibrium, neither can unilaterally adjust its parameters to further improve its objective without a corresponding change by the other. of the two-player game defined by the Equations (10) and (12), then $T^{*}(\bm{x},\bm{z})=\log q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})-\log p(\bm{z})$ (13) and $(\bm{\theta}^{*},\bm{\phi}^{*})$ is a global optimum of the variational lower bound in the Equation (8).
@@
Under misspecification such as inappropriate assumptions of latent variable distributions, AVB’s architecture can deliver a higher marginal likelihood theoretically. Meanwhile, by improving the marginal log-likelihood, one can reduce the reconstruction error $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]$, which measures the disparity between real and reconstructed data, and can be reduced by increasing marginal likelihood according to Equation (7). In the context of IRT, AVB thus stands out by capturing a more accurate respondents’ latent traits. Research in the computer science field also highlights that the combination of GAN and VAE can extract more discriminative information and provide more precise and clear model in the latent variable space (e.g., Makhzani et al.,, 2015; Hou et al.,, 2019).
@@ 3.1 Neural Network Design
Considering the schematic diagram in the Figure LABEL:fig:comp_avb, the encoder $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ and discriminator $T_{\bm{\psi}}(\bm{x},\bm{z})$ are designed to be multiple-layer feedforward neural networks (FNNs) in this study. A FNN is a type of artificial neural network where information flows in one direction, from the input layer to the output layer, through one or more hidden layers, without forming any cycles or loops. It is the simplest form of a neural network and by increasing the number of neurons in the hidden layer, any continuous function on a closed interval can be approximated at an arbitrary precision (Cybenko,, 1989).
FNN consists of layers of neurons, where each layer $l$ is represented mathematically by the following recursive relation: given the weight matrix $\bm{W}^{(l)}$ connecting layer l1l-1 and $l$, the $P_{l-1}\times 1$ output vector $\bm{h}^{(l-1)}$ from the previous layer l1l-1 and the $P_{l}\times 1$ bias vector $\bm{b}^{(l)}$ for layer $l$, the $P_{l}\times 1$ output $\bm{h}^{(l)}$ of the neurons in layer $l$ is defined by
@@
Parameter initialization is essential for training neural networks, as it sets initial values for weights and biases to ensure effective learning. Proper initialization prevents common training issues, such as vanishing or exploding gradients. In this study, Kaiming initialization (He et al.,, 2015) was chosen for the encoder and discriminator networks, as it is specifically designed for asymmetric activations such as GELU, maintaining stable variance across layers. For the decoder, viewed as a single-layer neural network with a sigmoid activation, Xavier initialization (Glorot and Bengio,, 2010) was applied to loading and intercept parameters to balance variance effectively.
Neural network parameters are commonly optimized using Stochastic Gradient Descent (SGD), an iterative method that updates parameters by minimizing a loss function computed from random batches of training data (LeCun et al.,, 2015). Formally, parameters $\bm{\omega}_{t}$ (e.g., $\bm{\theta},\bm{\phi},\bm{\psi}$ in the Algorithm 1) are updated according to $\bm{\omega}_{t+1}=\bm{\omega}_{t}-\eta\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega})$, where $\eta$ is the learning rate, and $\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega})$ is the gradient. Modern extensions of SGD methods, such as AdamW (Loshchilov,, 2017), improve on SGD by adaptive learning rates and decoupled weight decay. AdamW has become a popular choice in optimizing neural networks due to its ability to converge faster and improve model performance by effectively mitigating overfitting caused by large parameter values (Zhuang et al.,, 2022). Further details on initialization and AdamW optimization are provided in Appendix.
@@
In the Algorithm 1, the gradient descent step tries to force $T(\bm{x},\bm{z})$ close to $T^{*}(\bm{x},\bm{z})$ but the gap might fail to be sufficiently small. $T(\bm{x},\bm{z})$ tries to recover $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$, but this task can be quite challenging if there’s a large difference between $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the prior distribution $p(\bm{z})$. Therefore, Mescheder et al., (2017) also provided a technique to replace $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ with another KL divergence between $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and an adaptive distribution $r_{\alpha}(\bm{z}\mid\bm{x})$ rather than the simple, fixed Gaussian $p(\bm{z})$.
@@ 3.2 Adaptive Contrast Technique
To closely approximate the $q_{\phi}(\bm{z}\mid\bm{x})$, a tractable density choice of $r_{\alpha}(\bm{z}\mid\bm{x})$ is a Gaussian distribution with a diagonal covariance matrix, which matches the mean $\bm{\mu}(\bm{x})$ and variance $\bm{\sigma}^{2}(\bm{x})$ of $q_{\phi}(\bm{z}\mid\bm{x})$. Therefore, $r_{\alpha}(\bm{z}\mid\bm{x})$ is adaptive according to updating $\phi$. The KL term can also be approximated by the Theorem 2.1 and $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),r_{\alpha}(\bm{z}\mid\bm{x})\right]$ can be much smaller than $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$, which makes the construction of an effective discriminator easier (Mescheder et al.,, 2017). Moreover, given the distribution of the normalized vector $\tilde{\bm{z}}$, $\tilde{q}_{\bm{\phi}}(\tilde{\bm{z}}\mid\bm{x})$, reparameterization trick can further simplify the KL term into
$\text{KL}\left[\tilde{q}_{\bm{\phi}}(\tilde{\bm{z}}\mid\bm{x}),r_{0}(\tilde{\bm{z}})\right],\text{ where }r_{0}(\tilde{\bm{z}})\sim\mathcal{N}(\bm{0},\bm{I}),\tilde{\bm{z}}=\frac{\bm{z}-\bm{\mu}(\bm{x})}{\bm{\sigma}(\bm{x})}.$ (17)
@@ 3.3 Importance-weighted Technique
If the number of importance-weighted (IW) samples $R=1$, the IW-ELBO is reduced to the ELBO. As $R$ increases, IW-ELBO becomes more similar to the marginal likelihood than ELBO (Burda et al.,, 2015). Therefore, optimizing IW-ELBO, which is the importance-weighted Objective (8) or (16), achieves a better approximation to the MML estimator with slightly lower computational efficiency due to more samples but GPU computation source can ensure the computation time under control and even less than the time required by traditional CPU computation.
@@ 3.4 Implementation Summary
[…]
@@
    $\text{IW-ELBO}_{i}\leftarrow\frac{1}{S}\sum_{s=1}^{S}\left[\log\frac{1}{R}\sum_{r=1}^{R}w_{i,r,s}\right]$;
@@
 Optimization step: Perform SGD-updates via AdamW for $\bm{g}_{t}=(\bm{g}_{\bm{\theta}_{t}}^{T},\bm{g}_{\bm{\phi}_{t}}^{T},\bm{g}_{\bm{\psi}_{t}}^{T})$;
@@
Determining an optimal neural network configuration (e.g., the number and size of hidden layers) and other hyper-parameters, such as learning rates, are critical for satisfactory performance. GAN-based methods like AVB are particularly sensitive to these settings since they involve two interacting neural networks (Roth et al.,, 2017; Lucic et al.,, 2018). However, there is no panacea solution, as the architecture often depends on the specific application and the characteristics of the data. Empirical studies have shown that one or two hidden layers are typically sufficient to represent most continuous functions efficiently (Huang and Babri,, 1997; Huang,, 2003), and a maximum of $2n+1$ nodes can capture complex patterns (Hecht-Nielsen,, 1987). In the current study, various experimental settings were explored, and hyper-parameters that yielded the best results were selected. Experiments demonstrated that enhancing discriminator complexity (increased width or depth) generally improved feature learning, subsequently benefiting generator performance. However, an excessively complex discriminator negatively impacted generator learning, highlighting the necessity of careful hyperparameter tuning to maintain a proper balance.